OpenStudy (abbycross167):

Can someone please check my work on this math question? please? I'll fan and give a medal!!

OpenStudy (abbycross167):

The product of two consecutive odd integers is 63. Write a quadratic equation that you could solve to find the integers, then find the integers.

OpenStudy (abbycross167):

@JianEnriquez could you please try to help me?

jimthompson5910 (jim_thompson5910):

Let n be the first integer. The next integer after that is n+1. Since we want 2 odd integers, this means we need to jump to n+2 example: say n = 3, the next is n+1 = 4 and the next after that is n+2 = 5 So we're focused on n = 3 and n+2 = 5 In general, we have n and n+2 as our two unknown numbers. They multiply to get 63, so, (first number)*(second number) = 63 (n)*(n+2) = 63 n^2+2n = 63 Do you see what to do next?

OpenStudy (abbycross167):

(3)*(21+2) = 63 ?

jimthompson5910 (jim_thompson5910):

why not subtract 63 from both sides? n^2+2n = 63 n^2+2n-63 = 63-63 n^2+2n-63 = 0 Then use the quadratic formula

jimthompson5910 (jim_thompson5910):

Or you could factor, but factoring won't always work with quadratics. The quadratic formula always works with any quadratic.

OpenStudy (abbycross167):

oh okay so I just factored it and got (n−7)(n+9)

jimthompson5910 (jim_thompson5910):

good, so set that equal to zero and use the zero product property to solve for n (n-7)*(n+9) = 0 n-7 = 0 ... or ... n+9 = 0 n = ?? ... or ... n = ???

OpenStudy (abbycross167):

n=7 or n=−9 ?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

n = 7 n+2 = 7+2 = 9 so the two numbers are 7 and 9 7*9 = 63 -------------------------- n = -9 n+2 = -9+2 = -7 or you could say the two numbers are -9 and -7 -9*(-7) = 63

OpenStudy (abbycross167):

thank you very much!!! you're awesome!!

jimthompson5910 (jim_thompson5910):

glad to be of help